# Efficient Adversarial Training in LLMs with Continuous Attacks

### v2

@@ Abstract
Large language models (LLMs) are vulnerable to adversarial attacks that can bypass their safety guardrails. In many domains, adversarial training has proven to be one of the most promising methods to reliably improve robustness against such attacks. Yet, in the context of LLMs, current methods for adversarial training are hindered by the high computational costs required to perform discrete adversarial attacks at each training iteration. We address this problem by instead calculating adversarial attacks in the continuous embedding space of the LLM, which is orders of magnitudes more efficient. We propose a fast adversarial training algorithm (CAT) composed of two losses: the first makes the model robust on continuous embedding attacks computed on an adversarial behaviour dataset; the second ensures the usefulness of the final model by fine-tuning on utility data. Moreover, we introduce CAPO, an adversarial variant of IPO that does not require utility data for adversarially robust alignment. Our empirical evaluation on four models from different families (Gemma, Phi3, Mistral, Zephyr) and at different scales (2B, 3.8B, 7B) shows that both algorithms substantially enhance LLM robustness against discrete attacks (GCG, AutoDAN, PAIR), while maintaining utility. Our results demonstrate that robustness to continuous perturbations can extrapolate to discrete threat models. Thereby, we present a path toward scalable adversarial training algorithms for robustly aligning LLMs.
@@ 1 Introduction
We positively answer this research question using two novel adversarial training algorithms. We propose CAT, an efficient continuous AT algorithm, combining training on an adversarial behaviour dataset with fine-tuning on utility data. We further introduce CAPO, an adversarial variant of IPO that does not require utility data for adversarial alignment. We surpass the robustness-utility trade-offs of the discrete R2D2 AT algorithm [6], achieving up to $100\%$ attack robustness while requiring over $299$ times less computing resources. Additionally, we identify a failure mode in previous evaluation protocols: the models are tested with their chat template for safety evaluations but without it for utility evaluations. This protocol is unrealistic as the chat template is not enabled or disabled based on the prompt the user enters. By enabling the chat template for standard queries, we demonstrate that R2D2 overfits the safety objective and grammar of the harmful dataset. Thus, it often refuses to respond to benign inputs, thereby hurting its usefulness. In contrast, models trained with CAT and CAPO show substantially fewer refusals.
@@ 3 Method
In this section, we introduce our adversarial training (AT) algorithms: Continuous-Adversarial Training (CAT) and Continuous-Adversarial Preference Optimisation (CAPO). We begin by reviewing the standard AT regime from Madry et al. [4] (§ 3.1). We then explain differences between attacks in the standard AT setting and unique aspects of adversarial attacks in LLMs (§ 3.2). From there, we derive the Unlikelihood loss for—CAT (§ 3.3). Next, we introduce an adversarial IPO formulation—CAPO (§ 3.5). Finally, we discuss key design decisions in the above AT algorithm (§ 3.6).
@@ 3.1 Adversarial Training
AT is generally defined as a minimax optimization problem as follows [4]:
@@ 3.3 Adversarial Training in LLMs
3.4 Continuous-Adversarial Training
The primary difference between Mazeika et al. [6] and our method is the choice of perturbation set used during AT. Mazeika et al. [6] choose discrete suffix attacks $T_{\mathrm{suffix}}$ and employ the GCG algorithm along with several tricks to mitigate the computational cost to find a GCG attack. One optimisation they introduce is to only update the attack after every $k$ training steps. In contrast, we employ $T_{\mathrm{cont.}}$ with continuous attacks as introduced by Schwinn et al. [7], which are orders of magnitude (×$\times 299$) more efficient (see Table 1). Consequently, we do not require any additional tricks to further reduce computational costs. In the Unlikelihood loss (Eq 3) we add cut-off values for the toward and away loss to prevent over-optimising either. We implement this as $\mathcal{L}=\mathbb{I}[\mathcal{L}^{\prime}>c]0.999c+(\mathbb{I}[\mathcal{L}^{\prime}>c]0.001+\mathbb{I}[\mathcal{L}^{\prime}\leq c])\mathcal{L^{\prime}}$, where $c$ is the cutoff value chosen.
3.5 Continuous-Adversarial Preference Optimisation
Equation 3 has a similar form to DPO [30], which maximises the likelihood of a preferred answer while decreasing the likelihood of a dispreferred answer, given a prompt $x$. This motivates us to present the following loss function, which we will call Continuous-Adversarial Preference Optimisation (CAPO):
$\min_{\theta}-\mathbb{E}_{(x,y,\hat{y})\in\mathcal{D}}\left[\ell\left(\log\frac{f_{\theta}(y|x+\delta(x,\hat{y}))}{f_{\theta_{0}}(y|x)}-\log\,\frac{f_{\theta}(\hat{y}|x+\delta(x,\hat{y}))}{f_{\theta_{0}}(\hat{y}|x)}\right)\right],$ (5)
where $\ell(h)$ would be the $\log\sigma(h)$ in the original DPO, but we use the loss proposed in Azar et al. [31] called IPO, i.e. $\ell_{\beta}(h)=\left(h-\frac{1}{2\beta}\right)^{2}$, because it is less prone to overfitting. This loss implicitly minimises the Kullback-Leibler divergence w.r.t. the original model distribution $f_{\theta_{0}}(y|x)$, which prevents the model to collapse to degenerate behaviors leading to refuse all prompts with the refusal answer $y$. As a result, we are able to omit the utility dataset for CAPO.
@@ 4 Experimental Details
Moreover, we measure the utility of trained models using common benchmarks, including MMLU [35], Arc-E and Arc-C [36], and MT-Bench [37]. To reduce the computational demand, we evaluate $100$ questions for each category for MMLU. Finally, we introduce Harmless which consists of 40 harmless queries (e.g. Tell me a story, see App. G for full list) that are written in the same grammatical style as the Harmbench behaviour. We query the models with their chat template and report the number of refusals (checked manually). Note that only MT-Bench and Harmless use the model’s chat template.
@@
In our experiments, we adversarially fine-tuned four different open-source models Gemma [38], Phi-3-Mini [39], Mistral-7B [40], and Zephyr-7B [34] with increasing parameter counts—2B, 3.8B, 7B, 7B, respectively. We chose instruction-tuned models for all of them. We additionally include Zephyr + R2D2 in our evaluations, which is the Mistral-7B base model fine-tuned with the R2D2 AT algorithm [6]. This results in a diverse set of instruction-tuned models of different sizes. For more details, refer to App. A.2.
@@
We investigate two novel continuous AT algorithms in this work CAT and CAPO. Due to the computational complexity of fine-tuning LLMs, we do not perform full model fine-tuning for both methods but use LoRA [41] on all linear layers of the transformer architectures. Additionally, we use $4$-bit quantization for all training runs to further reduce the memory overhead. We use $\ell_{2}$ norm perturbations and set the size of the attack $\epsilon$ relatively to the average magnitude of the token embeddings of the respective model. For all models, we use $10$ attack iterations. We set $\epsilon=0.1$ for Gemma and Phi-3-Mini. For Mistral-7B and Zephyr-7B, we set $\epsilon=0.05$ and $\epsilon=0.075$, respectively. For a full list of AT hyperparameters, see App. A.1.
@@
We use three diverse adversarial attacks for the robustness evaluation. GCG, which has shown to achieve one of the highest average attack success rates (ASR) among other state-of-the-art attacks on several models [6]. Since GCG is a suffix attack, we further use AutoDAN and PAIR, which generate more diverse jailbreaks. Furthermore, PAIR has shown high ASR against previous AT approaches in LLMs [6]. To evaluate the ASR, we use the harmfulness classifier from [6], which was shown to align well with human judgement.
@@ 5 Results
Table 1: The combined number of forward (F) and backward (B) passes to compute a single adversarial example for different AT types. Further, the total number of F&B for the whole training and the number of training iterations and batch size are are shown.
@@
We first explore robustness extrapolation from continuous AT to discrete attacks for the CAT algorithm, which utilises additional utility data to maintain model performance. Figure 2 summarises the evaluation results. For all models, CAT considerably increases the average robustness against discrete adversarial attacks. For the Gemma and Zephyr models, robustness increases for all attacks. For Phi-3-Mini and Mistral-7B, PAIR still achieves high attack success rates (ASR). In terms of utility, we observe similar degradations for all CAT trained models. The MMLU and Arc scores decrease marginally, while the MT-Bench score decreases by approximately one. All models still show considerable utility after fine-tuning.
@@
(d) Zephyr-7B
Figure 2: Trade-off between utility and robustness for CAT (Eq. 4), CAPO (Eq. 5), and R2D2 [6], compared to their non-adversarially fine-tuned models. The objective is a small loss in utility and a large improvement in attack robustness. Larger is better for MMLU, Arc-E, Arc-C, MT-Bench (left of dashed line). Smaller is better for GCG, AutoDAN, and PAIR (right of dashed line). MT-Bench score is multiplied by 10 to see the change in performance on this $y$-axis. Error bars denote $95\%$ confidence intervals, illustrating the variation in ASR estimates based on the sample size of the datasets. CAPO is not provided for 7B models in (c, d), due to computational constraints.
@@
We further investigate if adversarial variations of proven alignment methods, such as IPO, can be used to align models in an adversarially robust manner (see Figure 2). For this purpose, we fine-tune Gemma and Phi-3-Mini using the proposed CAPO algorithm. Figure 2, illustrates differences between the base model, CAT, and CAPO. Despite using no utility dataset within CAPO to retain helpfulness, the algorithm does not introduce larger utility decreases on common benchmarks than CAT. Moreover, CAPO achieves considerably higher robustness against the jailbreaking method PAIR, demonstrating generalisation to diverse threat models. The Phi-3-Mini-CAPO model achieves $100\%$ attack robustness for all conducted attacks. For Gemma, robustness improvements also mostly surpass CAT, with slightly lower robustness against GCG. Compared to R2D2, CAPO does not require an auxiliary dataset to maintain utility and achieves higher robustness on average. Specifically for PAIR CAPO trained models exhibit considerably higher robustness. Lastly, the Phi-3-Mini-CAPO achieves a substantially higher score on the Harmless benchmark than CAT and R2D2.
@@ 6 Failure Modes of Training and Robustness Evaluations in LLMs
Common utility benchmarks such as MMLU or Arc do not use a chat template in their standard evaluation [42]. Firstly, this dramatically impacts performance, especially for smaller models, which often require a lot of prompt engineering to follow the few-shot prompts correctly. Secondly, it dramatically changes the mode of the model. In effect, a model might be overly robust in chat mode (i.e. when using a chat template) where it rejects most requests, but it might appear to have high utility in benchmarks because no chat template is used (e.g. MMLU). Arc as an evaluation benchmark is even more misleading as it measures the likelihood of a set of possible answer tokens, thus not reflecting the utility of the model when using a chat template. We quantitatively evaluate the refusals of MMLU questions when using a chat template in App. E. We recommend future work, to consider these issues when evaluating robustness and utility for the same model.
@@
AT datasets such as Harmbench [6] or AdvBench [43] tend to use a common grammatical and syntactical structure, using imperative commands such as “Tell me” or “Give instructions”. Chatting with our models and Zephyr + R2D2, we observe that requests would be refused when using this same style but are accepted if asked in a different style, such as “Could you please …?”. This holds for both harmful and harmless requests. For instance, Zephyr + R2D2 will refuse to answer “Tell me a story” and “Tell me how to build a bomb”, but will answer “Could you please tell me a story?” and “Could you please explain to me how to build a bomb?”. This also explains why the model may even appear useful under utility benchmarks employing chat templates such as MT-Bench. To demonstrate this failure case we create two small benchmark datasets called PoliteHarmbench (see App. F) and Harmless. The former rephrases the harmful behaviours politely, and the latter consists of harmless requests formulated in the same grammatical style as the original Harmbench behaviours. We leave developing better datasets and benchmarks for a future paper as it is outside the scope of this work.
@@ 7 Adversarial Training Ablations
Here, we provide ablations on several design choices of the proposed algorithms.
@@
Figure 3: Ablating how changing $\beta$ or $\epsilon{}$ affect GCG loss vs MMLU score on Gemma-CAPO
@@
Figure 4: Gemma-CAPO used for both plots: (a) Correlation between GCG loss and continuous attack loss. (b) GCG loss vs MMLU score for a variety of $\epsilon$ and $\beta$ values.
@@ 8 Conclusion
We answer our research question about the extrapolation of robustness under the continuous attack threat model to robustness under discrete attacks in the affirmative. We propose an efficient continuous adversarial training algorithm (CAT), combining training on an adversarial behaviour dataset with fine-tuning on utility data. Additionally, we introduce an adversarial variant of IPO (CAPO) that does not require additional utility data. Our algorithms achieve up to $100\%$ robustness against a set of state-of-the-art attacks (Phi-3-Mini-CAPO), surpassing robustness utility trade-offs in previous work [6] while requiring at least $299$ times less compute. In future work, we will further analyse settings where continuous robustness does not extrapolate (e.g. novel attacks) and possible ways to address this, such as larger and more diverse training data.
@@
We thank Maxime Darrin and Zichao Li for their helpful comments. This work is supported by CIFAR. This research was enabled in part by compute resources, software and technical help provided by Mila (mila.quebec).

### v3

@@ Abstract
Large language models (LLMs) are vulnerable to adversarial attacks that can bypass their safety guardrails. In many domains, adversarial training has proven to be one of the most promising methods to reliably improve robustness against such attacks. Yet, in the context of LLMs, current methods for adversarial training are hindered by the high computational costs required to perform discrete adversarial attacks at each training iteration. We address this problem by instead calculating adversarial attacks in the continuous embedding space of the LLM, which is orders of magnitudes more efficient. We propose a fast adversarial training algorithm (CAT) composed of two losses: the first makes the model robust on continuous embedding attacks computed on an adversarial behaviour dataset; the second ensures the usefulness of the final model by fine-tuning on utility data. Moreover, we introduce CAPO, an adversarial variant of IPO that does not require utility data for adversarially robust alignment. Our empirical evaluation on five models from different families (Gemma, Phi3, Mistral, Zephyr, Llama2) and at different scales (2B, 3.8B, 7B) shows that both algorithms substantially enhance LLM robustness against discrete attacks (GCG, AutoDAN, PAIR), while maintaining utility. Our results demonstrate that robustness to continuous perturbations can extrapolate to discrete threat models. Thereby, we present a path toward scalable adversarial training algorithms for robustly aligning LLMs.
@@ 1 Introduction
We positively answer this research question using two novel adversarial training algorithms. We propose CAT, an efficient continuous AT algorithm, combining training on an adversarial behaviour dataset with fine-tuning on utility data. We further introduce continuous adversarial preference optimisation (CAPO), an adversarial variant of identity preference optimisation (IPO) [11] that does not require utility data for adversarial alignment. We surpass the robustness-utility trade-offs of the discrete R2D2 AT algorithm [6], achieving up to $100\%$ attack robustness while requiring over $299$ times less computing resources. Additionally, we identify a failure mode in previous evaluation protocols: the models are tested with their chat template for safety evaluations but without it for utility evaluations. This protocol is unrealistic as the chat template is not enabled or disabled based on the prompt the user enters. By enabling the chat template for standard queries, we demonstrate that R2D2 overfits the safety objective and grammar of the harmful dataset. Thus, it often refuses to respond to benign inputs, thereby hurting its usefulness. In contrast, models trained with CAT and CAPO show substantially fewer refusals.
@@ 3 Method
In this section, we introduce our adversarial training (AT) algorithms: Continuous-Adversarial UL (CAT) and Continuous-Adversarial IPO (CAPO). We begin by reviewing the standard AT regime from Madry et al. [4] (§ 3.1). We then explain differences between attacks in the standard AT setting and unique aspects of adversarial attacks in LLMs (§ 3.2). From there, we derive the Unlikelihood loss for—CAT (§ 3.3). Next, we introduce an adversarial IPO formulation—CAPO (§ 3.5). Finally, we discuss key design decisions in the above AT algorithm (§ 3.6).
@@ 3.1 Adversarial Training
AT is generally defined as a minimax optimisation problem as follows [4]:
@@ 3.3 Adversarial Training in LLMs
3.4 Continuous-Adversarial Unlikelihood
The primary difference between Mazeika et al. [6] and our method is the choice of perturbation set used during AT. Mazeika et al. [6] choose discrete suffix attacks $T_{\mathrm{suffix}}$ and employ the GCG algorithm along with several tricks to mitigate the computational cost to find a GCG attack. One optimisation they introduce is to only update the attack after every $k$ training steps. In contrast, we employ $T_{\mathrm{cont.}}$ with continuous attacks as introduced by Schwinn et al. [7], which are orders of magnitude (×$\times 299$) more efficient (see Table 1). Consequently, we do not require any additional tricks to further reduce computational costs. In the Unlikelihood loss (Eq 3) we add cut-off values for the toward and away loss to prevent over-optimising either. Given a loss $\mathcal{L}^{\prime}$ before, we implement the cutoff as $\mathcal{L}=\mathbb{I}[\mathcal{L}^{\prime}>c]0.999c+(\mathbb{I}[\mathcal{L}^{\prime}>c]0.001+\mathbb{I}[\mathcal{L}^{\prime}\leq c])\mathcal{L^{\prime}}$, where $c$ is the cutoff value chosen.
3.5 Continuous-Adversarial IPO
Equation 3 has a similar form to DPO [31], which maximises the likelihood of a preferred answer while decreasing the likelihood of a dispreferred answer, given a prompt $x$. This motivates us to present the following loss function, which we will call Continuous-Adversarial IPO (CAPO):
$\min_{\theta}-\mathbb{E}_{(x,y,\hat{y})\in\mathcal{D}}\left[\ell_{\beta}\left(\log\frac{f_{\theta}(y|x+\delta(x,\hat{y}))}{f_{\theta_{0}}(y|x)}-\log\,\frac{f_{\theta}(\hat{y}|x+\delta(x,\hat{y}))}{f_{\theta_{0}}(\hat{y}|x)}\right)\right],$ (5)
where $\ell_{\beta}(h)$ would be the $\log\sigma(\beta h)$ in the original DPO, but we use the loss proposed in Azar et al. [11] called IPO, i.e. $\ell_{\beta}(h)=\left(h-\frac{1}{2\beta}\right)^{2}$, because it is less prone to overfitting. This loss implicitly minimises the Kullback-Leibler divergence w.r.t. the original model distribution $f_{\theta_{0}}(y|x)$, which prevents the model to collapse to degenerate behaviors leading to refuse all prompts with the refusal answer $y$. As a result, we are able to omit the utility dataset for CAPO.
@@ 4 Experimental Details
Moreover, we measure the utility of trained models using common benchmarks, including MMLU [35], Arc-E and Arc-C [36], and MT-Bench [37]. To reduce the computational demand, we evaluate $100$ questions for each category for MMLU. Finally, we introduce Harmless which consists of 40 harmless queries (e.g. Tell me a story, see App. I for full list) that are written in the same grammatical style as the Harmbench behaviour. We query the models with their chat template and report the number of refusals (checked manually). Note that only MT-Bench and Harmless use the model’s chat template.
@@
In our experiments, we adversarially fine-tuned four different open-source models Gemma [38], Phi-3-Mini [39], Mistral-7B [40], Zephyr-7B [34], and Llama2-7B [41] with increasing parameter counts—2B, 3.8B, 7B, 7B, and 7B, respectively. We chose instruction-tuned models for all of them. We additionally include Zephyr + R2D2 in our evaluations, which is the Mistral-7B base model fine-tuned with the R2D2 AT algorithm [6]. This results in a diverse set of instruction-tuned models of different sizes. For more details, refer to App. A.2.
@@
We investigate two novel continuous AT algorithms in this work CAT and CAPO. Due to the computational complexity of fine-tuning LLMs, we do not perform full model fine-tuning for both methods but use LoRA [42] on all linear layers of the transformer architectures. Additionally, we use $4$-bit quantization for all training runs to further reduce the memory overhead. We use $\ell_{2}$ norm perturbations and set the size of the attack $\epsilon$ relatively to the average magnitude of the token embeddings of the respective model. For all models, we use $10$ attack iterations. We set $\epsilon=0.1$ for Gemma and Phi-3-Mini. For Mistral-7B, Llama-7B, and Zephyr-7B, we set $\epsilon=0.05$, $\epsilon=0.05$, and $\epsilon=0.075$, respectively. For a full list of AT hyperparameters, see App. A.1.
@@
We use three diverse adversarial attacks for the robustness evaluation. GCG, which has shown to achieve one of the highest average attack success rates (ASR) among other state-of-the-art attacks on several models [6]. Since GCG is a suffix attack, we further use AutoDAN and PAIR, which generate more diverse jailbreaks. Finally, we also evaluate against Adaptive Attacks [2] and ICL [19] (see Table 5 and Table 6). Furthermore, PAIR has shown high ASR against previous AT approaches in LLMs [6]. To evaluate the ASR, we use the harmfulness classifier from [6], which was shown to align well with human judgement.
@@ 5 Results
Table 1: The combined number of forward (F) and backward (B) passes to compute a single adversarial example for different AT types. The total number of F&B for the whole training and the number of training iterations and batch size are are shown. Time is the wallclock time for a single batch weight update (measured on 1 A100 with Mistral).
@@
We first explore robustness extrapolation from continuous AT to discrete attacks for the CAT algorithm, which utilises additional utility data to maintain model performance. Figure 2 summarises the evaluation results. For all models, CAT considerably increases the average robustness against discrete adversarial attacks. For the Gemma and Zephyr models, robustness increases for all attacks. For Phi-3-Mini and Mistral-7B, PAIR still achieves high attack success rates (ASR). In terms of utility, we observe similar degradations for all CAT trained models. All models still show considerable utility after fine-tuning.
@@
(d) Llama-7B
(e) Zephyr-7B
Figure 2: Trade-off between utility and robustness for CAT (Eq. 4), CAPO (Eq. 5), and R2D2 [6], compared to their non-adversarially fine-tuned models. The objective is a small loss in utility and a large improvement in attack robustness. Larger is better for MMLU, Arc-E, Arc-C, MT-Bench (left of dashed line). Smaller is better for GCG, AutoDAN, and PAIR (right of dashed line). MT-Bench score is multiplied by 10 to see the change in performance on this $y$-axis. Additional results are included in App. B.
@@
We further investigate if adversarial variations of proven alignment methods, such as IPO, can be used to align models in an adversarially robust manner (see Figure 2). For this purpose, we fine-tune Gemma and Phi-3-Mini using the proposed CAPO algorithm. Figure 2, illustrates differences between the base model, CAT, and CAPO. Despite using no utility dataset within CAPO to retain helpfulness, the algorithm does not introduce larger utility decreases on common benchmarks than CAT. Moreover, CAPO achieves considerably higher robustness against the jailbreaking method PAIR, demonstrating generalisation to diverse threat models. The Phi-3-Mini-IPO model achieves $100\%$ attack robustness for all conducted attacks. For Gemma, robustness improvements also mostly surpass CAT, with slightly lower robustness against GCG. Compared to R2D2, CAPO does not require an auxiliary dataset to maintain utility and achieves higher robustness on average. Specifically for PAIR CAPO trained models exhibit considerably higher robustness. Lastly, the Phi-3-Mini-IPO achieves a substantially higher score on the Harmless benchmark than CAT and R2D2.
@@ 6 Failure Modes of Training and Robustness Evaluations in LLMs
Common utility benchmarks such as MMLU or Arc do not use a chat template in their standard evaluation [43]. Firstly, this dramatically impacts performance, especially for smaller models, which often require a lot of prompt engineering to follow the few-shot prompts correctly. Secondly, it dramatically changes the mode of the model. In effect, a model might be overly robust in chat mode (i.e. when using a chat template) where it rejects most requests, but it might appear to have high utility in benchmarks because no chat template is used (e.g. MMLU). Arc as an evaluation benchmark is even more misleading as it measures the likelihood of a set of possible answer tokens, thus not reflecting the utility of the model when using a chat template. We quantitatively evaluate the refusals of MMLU questions when using a chat template in App. G. We recommend future work, to consider these issues when evaluating robustness and utility for the same model.
@@
AT datasets such as Harmbench [6] or AdvBench [44] tend to use a common grammatical and syntactical structure, using imperative commands such as “Tell me” or “Give instructions”. Chatting with our models and Zephyr + R2D2, we observe that requests would be refused when using this same style but are accepted if asked in a different style, such as “Could you please …?”. This holds for both harmful and harmless requests. For instance, Zephyr + R2D2 will refuse to answer “Tell me a story” and “Tell me how to build a bomb”, but will answer “Could you please tell me a story?” and “Could you please explain to me how to build a bomb?”. This also explains why the model may even appear useful under utility benchmarks employing chat templates such as MT-Bench. To demonstrate this failure case we create two small benchmark datasets called PoliteHarmbench (see App. H) and Harmless. The former rephrases the harmful behaviours politely, and the latter consists of harmless requests formulated in the same grammatical style as the original Harmbench behaviours. We leave developing better datasets and benchmarks for a future paper as it is outside the scope of this work.
@@ 7 Adversarial Training Ablations
Figure 3: Ablating how changing $\beta$ or $\epsilon{}$ affect GCG loss vs MMLU score on Gemma-IPO
@@
Figure 4: Gemma-IPO used for both plots: (a) Correlation between GCG loss and continuous attack loss. (b) GCG loss vs MMLU score for a variety of $\epsilon$ and $\beta$ values.
@@ 8 Conclusion
We answer our research question about the extrapolation of robustness under the continuous attack threat model to robustness under discrete attacks in the affirmative. We propose an efficient continuous adversarial training algorithm (CAT), combining training on an adversarial behaviour dataset with fine-tuning on utility data. Additionally, we introduce an adversarial variant of IPO (CAPO) that does not require additional utility data. Our algorithms achieve up to $100\%$ robustness against a set of state-of-the-art attacks (Phi-3-Mini-CAPO), surpassing robustness utility trade-offs in previous work [6] while requiring at least $299$ times less compute. In future work, we will further analyse settings where continuous robustness does not extrapolate (e.g. novel attacks) and possible ways to address this, such as larger and more diverse training data. Additionally, the objectives of preventing harmful output and machine unlearning are closely related as such the applicability of our method for machine unlearning would be an interesting angle for further exploration.
@@
We thank Maxime Darrin, Zichao Li, and the anonymous reviewers for their helpful comments. We thank Mato Gudelj for code in running the NPO baseline. This work is supported by CIFAR. This research was enabled in part by compute resources, software and technical help provided by Mila (mila.quebec). Leo Schwinn gratefully acknowledges funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) - Projectnumber 544579844. Leo Schwinn acknowledges travel support from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 951847.